"In the triangle ABC, measure of angle A= 60 degrees and the measure of angel C = 80 degrees. The exterior angle at B measures(x+10).
What is the value of x?"
Show all work!!!

Answer by stanbon(48535)

Angle B = 180-(60+80)=40 degrees
The exterior angle at B is 180-40=140 degrees
x+10=140
x=130
Cheers,
Stan H.